MONT 102N – Modeling the Environment Paper 2 – Human Effects on the Natural World in Massachusetts and Montana Due Date: Friday November 4

General Information

The second larger paper assignment for the semester will involve tying together:

• A chapter in Collapse that we did not discuss very much or at all–the Montana chapter at the start (Chapter 1),

• Some of the history of land use in Massachusetts that we will see on the Harvard Forest trips September 17 and October 22, and

• Some of the scientific work carried out at Harvard Forest on evolution of landscapes over time with human interaction, detailed in the articles posted on the course homepage. Your essay should be about 5 double-spaced pages in length (but don’t worry if you go

a bit over or are a bit short – the length is not the main point). Prepare your essays using Microsoft Word or equivalent software, and submit your work to me by email to:

                           jlittle@holycross.edu

This is due no later than 5:00pm on Friday, November 4.

The General Topic

Consider time frames from the immediate decades before European settlement, to the century or two after European settlement, to the present in both Montana and Massachusetts. How have the landscape and environments (especially the plant and animal species present) in these two areas changed over time? How do we know that – how is it possible to quantify and measure these changes? (You will have more information here about the situation in Massachusetts from observations you make on the field trips and from the two articles posted on the course homepage. You can use that information without trying to say anything parallel about Montana because many of the same techniques would apply to the study of other geographical areas too.) Are there similarities in the two courses of development? Are there differences? Have people’s cultural values influenced their choices about land use and possibly contributed to environmental problems in these areas? Is Montana similar to the Northeast in terms of suitability for agriculture? Would the landscape of Massachusetts look different now if the U.S. ended at the western borders of Vermont, Massachusetts and Connecticut (if the U.S. had not expanded to the west)?

More Details

This will probably be a bit more of a “research paper” than the first assignment. You will need to consult at least Collapse and the articles mentioned above. The “they say/I say” framework will still be useful to keep in mind. A good paper should definitely have a specific point of view about the questions above, and the first thing you will want to do is to identify what it is you want to say and how you will tie everything together.

As always, write in a reasonably formal style incorporating correct grammar, spelling, and usage. Your essay should be well-structured with paragraphs corresponding to the major points in your outline, arranged in a logical sequence. (Feel free to use some of the “templates” from “They Say/I Say” to help structure things – no footnotes are necessary for that.) Your job is to convince your reader of the soundness of your conclusions, starting “from the beginning.” That is, you should not assume the reader has been present for any of the discussions that we have had in class, or that you might have had with other Natural World Cluster students. You might try explaining possible opposing points of view (“introducing a naysayer”), then giving your reasons for rejecting them, as part of your argument.

I will expect to see specific details from your sources cited as evidence to support your assertions, and in many cases short quotations from the other text will be the most effective way to provide evidence. Please identify any such quotations by the page(s) on which they appear. Use the following style:

Diamond writes that our world “cannot sustain China and other Third World countries and current First World countries all operating at First World levels.” (Diamond, p. 376).

Use of other sources and direct quotations or other information should follow the following guidelines:

1) Identify direct quotations with quotation marks

2) All direct quotations and paraphrases of arguments found in Collapse or other sources should be identified by notes in parentheses giving information where the quote can be found within your source,

3) List all your sources in a References section at the end of your paper:

a) For books, give the author(s), the title, city where the publisher is located, the publisher, and year of publication.

b) For magazine or newspaper articles, give the author(s), the title of the article, the title of the magazine or newspaper, the date of publication, and the starting and final page number of the specific article.

c) For web sites, give an author (if you can determine that), the full URL (web address) and the date you accessed it (since information on the web often changes!)